OpenStudy (anonymous):

If anyone is really good with proofs please let me know. I have a couple questions concerning proofs and I don't understand them at all.

OpenStudy (anonymous):

well, if you can show me, i'd be willing to try...

OpenStudy (compassionate):

Ask it.

OpenStudy (anonymous):

Given: Square with side c. All four interior triangles are right triangles. All four interior triangles are congruent. The interior quadrilateral is a square. Prove: a2 + b2 = c2 When written in the correct order, the paragraph below proves the Pythagorean Theorem using the diagram. Let a represent the height and b represent the base of each triangle. The area of one triangle is represented by the expression ab. (1) The length of a side of the interior square is (a – b). (2) The area of all four triangles will be represented by 4 • ab or 2ab. (3) The area of the interior square is (a – b)2. (4) By distribution, the area is a2 – 2ab + b2. The area of the exterior square is found by squaring side c, which is c2, or by adding the areas of the four interior triangles and interior square, 2ab + a2 – 2ab + b2. Therefore, c2 = 2ab + a2 – 2ab + b2. Through addition, c2 = a2 + b2. Which is the most logical order of statements (1), (2), (3), (4) to complete the proof?

OpenStudy (anonymous):

If you need the picture, I can link that too.

OpenStudy (anonymous):

I figure the most logical order would be (1),(2),(4),(3)

OpenStudy (shayaan_mustafa):

Show diagrams.